To find an equation of a line given the slope and one point on the line
m = 4; (-1,1)
2

Answer by rchill(405)

Substitute your values into the equation y=mx+b, so you'll have 1=4*(-1)+b, which reduces to b=5. That means the equation of your line is y=4x+5.